the potential difference of a cell on open circuit is 6V which falls to 4V when a current of 4A is drawn from the cell.find the internal resistance of the cell.
1
Expert's answer
2012-06-14T08:06:06-0400
R = (V1 - V0) / I = (6 - 4) / 4 = 1 Ohm